Wildhorse Candy Company offers a coffee mug as a premium for every ten $1 candy bar wrappers presented by customers together with $2. The purchase price of each mug to the company is $2.40; in addition it costs $1.60 to mail each mug. The results of the premium plan for the years 2020 and 2021 are as follows (assume all purchases and sales are for cash):
2020 2021
Coffee mugs purchased 750,000 780,000
Candy bars sold 5,500,000 6,690,000
Wrappers redeemed 2,600,000 4,230,000
2020 wrappers expected to be redeemed in 2021 2,000,000
2021 wrappers expected to be redeemed in 2022 2,710,000

Prepare the general journal entries that should be made in 2020 and 2021 related to the above plan by Wildhorse Candy.

Indicate the account names, amounts, and classifications of the items related to the premium plan that would appear on the Wildhorse Candy Company balance sheet and income statement at the end of 2020 and 2021.

Balance Sheet

Account Name Class 2020 2021

Operating IncomeCurrent LiabilityOperating ExpenseCurrent Asset

$

$

Operating IncomeCurrent AssetCurrent LiabilityOperating Expense

$

$


Income Statement
Account Name Class 2020 2021

    Operating Expense    Operating Income    Current Asset    Current Liability    

$

$
Answer #1

1.

No. General Journal Debit Credit
2020
1 Inventory of Premiums (750000 x $2.40) 1800000
Cash 1800000
(To record the purchase of premium inventory)
2 Cash 5500000
Sales Revenue (5500000 x $1) 5500000
(To record the sales)
3 Cash [2600000/10 = 260000 x ($2.00 - $1.60)] 104000
Premium Expense [260000 x ($2.40 + $1.60 - $2.00)] 520000
Inventory of Premiums (260000 x $2.40) 624000
(To record the wrappers redeemed)
4 Premium Expense 400000
Premium Liability [2000000/10 = 200000 x ($2.40 + $1.60 - $2.00)] 400000
(To record the premium liability)
2021
1 Inventory of Premiums (780000 x $2.40) 1872000
Cash 1872000
(To record the purchase of premium inventory)
2 Cash 6690000
Sales Revenue (6690000 x $1) 6690000
(To record the sales)
3 Cash [4230000/10 = 423000 x ($2.00 - $1.60)] 169200
Premium Liability 400000
Premium Expense [(423000 - 200000) x ($2.40 + $1.60 - $2.00)] 446000
Inventory of Premiums (423000 x $2.40) 1015200
(To record the wrappers redeemed)
4 Premium Expense 542000
Premium Liability [2710000/10 = 271000 x ($2.40 + $1.60 - $2.00)] 542000
(To record the premium liability)

2.

Balance Sheet
Account Name Class 2020 2021
Inventory of Premiums Current Asset $       1,176,000 $       2,032,800
Premium Liability Current Liability $         400,000 $         542,000
Income Statement
Account Name Class 2020 2021
Premium Expense Operating Expense $         920,000 $         988,000